# MILR: Improving Multimodal Image Generation via Test-Time Latent Reasoning

### v1

@@ 4.3.1 Hyperparameters
Hu et al. (2025) have empirically shown that, in autoregressive image generation, early-stage tokens govern the overall image structure, and perturbing the first 20% of image tokens results in significant structural deviations. Therefore, we conservatively adopt a very small $\lambda_{v}$, varying it from 0.01 to 0.1 (see Figure 5 (b)). Surprisingly, optimizing only the first 2% of all image tokens already yields peak performance, while further increasing $\lambda_{v}$ tends to degrade it (see examples in Figure 12).
@@ 4.4 Reward Models
The reward model is a crucial component of MILR, providing learning signals for latent reasoning. Following previous work (Liu et al., 2025a; Tong et al., 2025), we have used the benchmark’s evaluator as the reward model (denoted by [Uncaptioned image] OracleReward), but in real-world scenarios, oracle rewards are usually unknown, and it is difficult to design domain-specific rewards. To show that MILR is effective without the reliance on [Uncaptioned image] , we test it with a set of off-the-shelf reward models on GenEval:
@@
  • [Uncaptioned image] GPT-4o represents a frontier critic for assessing image quality (Hurst et al., 2024).
  • [Uncaptioned image] UnifiedReward is specifically tuned for a unified evaluation of MUG (Wang et al., 2025).
  • [Uncaptioned image] MixedReward is a composite critic for more comprehensive evaluation. It aggregates rewards from specialized models, including GroundingDINO (Liu et al., 2024) (evaluating object detection), GIT (Wang et al., 2022) (judging colors), and [Uncaptioned image] (assessing aesthetics).
@@ 5 Discussion
We acknowledge two primary limitations of MILR. First, we have focused on an implementation of MILR that builds upon autoregressive MUG, where both text and image tokens are generated autoregressively. Another strong paradigm of MUG is through diffusion image generation, where image tokens within the same image can attend to each other (Deng et al., 2025). It is interesting to see if the effectiveness of MILR transfers. Second, MILR relies on a reward model for learning signals, but, in practice, a perfect reward model usually does not exist, and it is difficult to design a domain-agnostic reward model. In addition, our experimental results have revealed that the strongest non-oracle reward model still lags behind the oracle reward. Thus, future work is well-suited for designing reward models that can generalize like the unified reward model of Wang et al. (2025).
@@ 6 Conclusion
Reproducibility statement
The code will be released soon. In Section 4.1, we also detailed the hyperparameters used in our experiments.
@@ A.3.2 Analysis of Reward models
In Section 4.4, we evaluated MILR with different reward models: [Uncaptioned image] SelfReward, [Uncaptioned image] GPT-4o, [Uncaptioned image] UnifiedReward, [Uncaptioned image] MixedReward, and [Uncaptioned image] OracleReward. We further provide example images generated under each setup in Figure 13. Overall, [Uncaptioned image] establishes an upper bar of generation quality, and [Uncaptioned image] , among non-oracle reward models, performs best because it incorporates multiple specialized critics to provide a comprehensive assessment. Though [Uncaptioned image] still lags behind [Uncaptioned image] , in the absence of [Uncaptioned image] , we can compose a strong universal reward model from specialized off-the-shelf image critics.
@@ A.3.3 Efficieny Analysis
We summarize the computing resources used by different image generation models in Table 7. Compared with training-based reasoning approaches such as T2I-R1 (Jiang et al., 2025a), Janus-Pro with DPO/GRPO (Tong et al., 2025), and Flow-GRPO (Liu et al., 2025a), our test-time reasoning method achieves the best performance, without relying on curated reasoning data for training. As an example, Flow-GRPO requires about 2K A800 GPU hours for training but only matches MILR, not to mention its labor cost in curating reasoning data. Moreover, when compared with the strong test-time reasoning method Best-of-N (N=20 vs. T=20 in MILR), MILR surpasses it by 0.04 while requiring less inference time because it employs an early-stop strategy, i.e., it stops once the generated image satisfies evaluation criteria.
@@
T2I-R1 (Jiang et al., 2025a) 8 H800 16 h 0.79
Janus-Pro+GRPO (Tong et al., 2025) 8 A100 $\sim$9 h 0.81
Janus-Pro+DPO (Tong et al., 2025) 8 A100 $\sim$9 h 0.83
Flow-GRPO (Liu et al., 2025a) 24 A800 $\sim$100 h 0.95
Reflect-DiT (Li et al., 2025b) A6000 24 h 16 h 0.81

### v2

@@ 4.3.1 Hyperparameters
hu2025anchortokenmatchingimplicit have empirically shown that, in autoregressive image generation, early-stage tokens govern the overall image structure, and perturbing the first 20% of image tokens results in significant structural deviations. Therefore, we conservatively adopt a very small $\lambda_{v}$, varying it from 0.01 to 0.1 (see Figure 5 (b)). Surprisingly, optimizing only the first 2% of all image tokens already yields peak performance, while further increasing $\lambda_{v}$ tends to degrade it (see examples in Figure 12).
@@ 4.4 Reward Models
The reward model is a crucial component of MILR, providing learning signals for latent reasoning. Following previous work (liu2025flowgrpo; tong2025delving), we have used the benchmark’s evaluator as the reward model (denoted by [Uncaptioned image] OracleReward), but in real-world scenarios, oracle rewards are usually unknown, and it is difficult to design domain-specific rewards. To show that MILR is effective without the reliance on [Uncaptioned image] , we test it with a set of off-the-shelf reward models on GenEval:
@@
  • [Uncaptioned image] GPT-4o represents a frontier critic for assessing image quality (hurst2024gpt4o).
  • [Uncaptioned image] UnifiedReward is specifically tuned for a unified evaluation of MUG (wang2025unifiedreward).
  • [Uncaptioned image] MixedReward is a composite critic for more comprehensive evaluation. It aggregates rewards from specialized models, including GroundingDINO (liu2024grounding) (evaluating object detection), GIT (wang2022git) (judging colors), and [Uncaptioned image] (assessing aesthetics).
@@ 5 Discussion
We acknowledge two primary limitations of MILR. First, we have focused on an implementation of MILR that builds upon autoregressive MUG, where both text and image tokens are generated autoregressively. Another strong paradigm of MUG is through diffusion image generation, where image tokens within the same image can attend to each other (deng2025bagel). It is interesting to see if the effectiveness of MILR transfers. Second, MILR relies on a reward model for learning signals, but, in practice, a perfect reward model usually does not exist, and it is difficult to design a domain-agnostic reward model. In addition, our experimental results have revealed that the strongest non-oracle reward model still lags behind the oracle reward. Thus, future work is well-suited for designing reward models that can generalize like the unified reward model of wang2025unifiedreward.
@@ A.3.2 Analysis of Reward models
In Section 4.4, we evaluated MILR with different reward models: [Uncaptioned image] SelfReward, [Uncaptioned image] GPT-4o, [Uncaptioned image] UnifiedReward, [Uncaptioned image] MixedReward, and [Uncaptioned image] OracleReward. Table 6, Table 7 and Table 8 show results across benchmarks with different rewards. We further provide example images generated under each setup in Figure 13. Overall, [Uncaptioned image] establishes an upper bar of generation quality, and [Uncaptioned image] , among non-oracle reward models, performs best because it incorporates multiple specialized critics to provide a comprehensive assessment. Though [Uncaptioned image] still lags behind [Uncaptioned image] , in the absence of [Uncaptioned image] , we can compose a strong universal reward model from specialized off-the-shelf image critics.
@@
Table 7: T2I-CompBench results with different reward models. The best performance is in bold.
Reward Color Shape Texture Spatial Non-Spatial Complex Overall Baseline 0.6359 0.3528 0.4936 0.2061 0.3085 0.3559 0.3921 [Uncaptioned image] Self Reward 0.7196 0.4620 0.5887 0.2379 0.3055 0.3868 0.4501 [Uncaptioned image] GPT-4o 0.7624 0.4701 0.6561 0.2778 0.3102 0.3881 0.4775 [Uncaptioned image] UnifiedReward 0.8208 0.4609 0.5835 0.3210 0.3044 0.3700 0.4651 [Uncaptioned image] MixedReward 0.8009 0.4765 0.6608 0.4077 0.3055 0.3745 0.5043 [Uncaptioned image] OracleReward 0.8508 0.5117 0.6949 0.4613 0.3078 0.3684 0.5325
Table 8: WISE results with different reward models. The best performance is in bold.
Reward Cultural Time Space Biology Physics Chemistry Overall Base 0.30 0.37 0.49 0.36 0.42 0.26 0.35 [Uncaptioned image] Self Reward 0.40 0.43 0.53 0.38 0.44 0.23 0.41 [Uncaptioned image] GPT-4o 0.53 0.53 0.62 0.52 0.53 0.34 0.52 [Uncaptioned image] UnifiedReward 0.45 0.54 0.60 0.46 0.52 0.30 0.48 [Uncaptioned image] MixedReward 0.48 0.51 0.61 0.44 0.51 0.33 0.49 [Uncaptioned image] Metric Reward 0.64 0.65 0.72 0.66 0.71 0.37 0.63
@@ A.3.3 Efficieny Analysis
We summarize the computing resources used by different image generation models in Table 9. Compared with training-based reasoning approaches such as T2I-R1 (jiang2025t2i-r1), Janus-Pro with DPO/GRPO (tong2025delving), and Flow-GRPO (liu2025flowgrpo), our test-time reasoning method achieves the best performance, without relying on curated reasoning data for training. As an example, Flow-GRPO requires about 2K A800 GPU hours for training but only matches MILR, not to mention its labor cost in curating reasoning data. Moreover, when compared with the strong test-time reasoning method Best-of-N (N=20 vs. T=20 in MILR), MILR surpasses it by 0.04 while requiring less inference time because it employs an early-stop strategy, i.e., it stops once the generated image satisfies evaluation criteria.
@@
T2I-R1 (jiang2025t2i-r1) 8 H800 16 h 0.79
Janus-Pro+GRPO (tong2025delving) 8 A100 $\sim$9 h 0.81
Janus-Pro+DPO (tong2025delving) 8 A100 $\sim$9 h 0.83
Flow-GRPO (liu2025flowgrpo) 24 A800 $\sim$100 h 0.95
Reflect-DiT (li2025reflectdit) A6000 24 h 16 h 0.81